OpenStudy (anonymous):

Differentiate: f(theta)=ln(2theta) with respect to theta (I'm learning derivatives of logarithmic functions)

OpenStudy (agent0smith):

\[\Large f( \theta )=\ln(2 \theta ) \] The derivative of ln[f(x)] is \[\Large \frac{ d }{ dx } \ln [f(x)] = f'(x)* \frac{ 1 }{ f(x) }\] Derivative of 2theta is just two, so: \[\large \frac{ d }{ dx } \ln(2 \theta ) = 2 * \frac{ 1 }{ 2 \theta }\]

OpenStudy (agent0smith):

ie, differentiate the inside function (inside the log), and multiply by 1 over the inside function.

OpenStudy (anonymous):

so i had a question about derivatives of logarithmic functions in calc.

OpenStudy (anonymous):

is it basically inverse? if its ln?

OpenStudy (anonymous):

I was watching the lecture but I wasn't sure what exactly it was about..

OpenStudy (agent0smith):

Yeah kinda.... the simplest way i can think of is like i said above, differentiate the inside function (inside the log), and multiply that by 1 over the inside function.

OpenStudy (anonymous):

is that my final answer?

OpenStudy (anonymous):

or would it be 1/theta because the two's cancel?

OpenStudy (agent0smith):

eg \[\large \frac{ d }{ dx} \ln(x^2) = 2x * \frac{ 1 }{ x^2 }\]

OpenStudy (anonymous):

and whatever your writing... please continue lol don't mean to interrupt

OpenStudy (agent0smith):

1/theta because the two's cancel.

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

what about this one:

OpenStudy (anonymous):

f(t) = ln [sqrt(t^3)-6]

OpenStudy (agent0smith):

Derivative of the inside function is (change the sqrt to an exponent to simplfify first) \[\Large \frac{ d }{dt} ((t^3)^{1/2}-6 ) = \Large \frac{ d }{dt} (t^{3/2}-6 ) = \frac{ 3 }{ 2 }t^{1/2}\]

OpenStudy (agent0smith):

Then multiply that by 1/[sqrt(t^3)-6]

OpenStudy (anonymous):

is that 3t^2 / (2t^3 - 12)?

OpenStudy (agent0smith):

\[\Large f \prime (t) = \frac{ 3 }{ 2 }t^{1/2} * \frac{ 1 }{ \sqrt{t^3}-6 } = \] if you meant this, then yes \[\Large f \prime (t) = \frac{ 3 t^{1/2} }{ 2\sqrt{t^3}-12 }\]

OpenStudy (anonymous):

hmm... its slightly different then what everyone else got.

OpenStudy (agent0smith):

Well you can write it a little differently, with the exponents and all.

OpenStudy (anonymous):

okthanks!

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=derivative+ln+%5Bsqrt%28t%5E3%29-6%5D+ if you scroll down, you'll see my solution (they just used the sqrt sign instead of a 1/2 exponent).

OpenStudy (anonymous):

ok will do.